# On long-duration storage, weather uncertainty and limited foresight

### v1

@@ Abstract
Long-duration energy storage (LDES) is a key component for fully renewable, sector-coupled energy systems based on wind and solar. While capacity expansion planning has begun to take into account interannual weather variability, it often ignores weather uncertainty and limited foresight in capacity and operational decisions. We build a stochastic capacity expansion model for fully decarbonized energy systems with LDES in Europe accounting for weather uncertainty - isolating the effect of limited foresight by comparing it to a perfect foresight benchmark. Under limited foresight, LDES acts as a hedge against extreme system states operating defensively and exhibiting a stockpiling effect absent under perfect foresight. Solar PV gains in system value for its higher predictability with up to 29% higher capacities versus the benchmark while onshore wind capacities are lower. We shed light on the underlying mechanisms by deriving implicit LDES bidding curves. We show that LDES bids reflect the costs and the weather-dependent probability of extreme system states conditional on the current system state. This has important implications for the price formation on renewable electricity markets, as a wide and continuous range of probabilistic LDES bids alleviates concerns of extreme price disparity at high renewable shares.
@@ 1 Introduction
The energy sector is at the center of climate change mitigation efforts around the world [international_energy_agency_net_2024, intergovernmental_panel_on_climate_change_ipcc_technical_2023]. Variable renewable energy (VRE) including wind and solar energy are likely to supply the majority of the energy in a decarbonized system for their ubiquitous availability, cost competitiveness and ease of deployment [schlachtberger_cost_2018, victoria_speed_2022, brown_response_2018, breyer_history_2022, luderer_impact_2022, hansen_status_2019]. The resulting system dependency on weather variables such as wind speeds, solar irradiation, water inflows and temperatures due to electrified heating and cooling demand creates flexibility requirements on various temporal and spatial scales [lund_review_2015, levin_energy_2023].
@@
Long-duration energy storage (LDES) is considered a key enabler of very high VRE shares in a decarbonized future by providing seasonal or interannual balancing and bridging periods of low solar and wind feed-in that may last multiple days or even weeks, so-called Dunkelflauten [lopez_prol_economics_2021, dowling_role_2020, ruhnau_storage_2022, sepulveda_design_2021, kittel_coping_2025, staadecker_value_2024]. LDES is not well defined but often used to refer to storage technologies that offer durations between ten and hundreds of hours and which are characterized by low energy capacity costs [dowling_long-duration_2021, albertus_long-duration_2020, brown_ultra-long-duration_2023]. While several technology options exist, most of the academic literature, planned legislation and current market developments in Europe and the US suggest that hydrogen storage in salt caverns in conjunction with electrolysis and hydrogen gas turbines (power-to-gas-to-power) will play a dominant role [dowling_role_2020, federal_ministry_for_economic_affairs_and_climate_action_of_germany_weisbuch_2025, rwe_erweiterungsprojekt_nodate, uniper_press_nodate].
@@
Capacity expansion models (CEM) are well established and important tools for policymakers, investors and other stakeholders in the energy transition. Yet, the representation of LDES in these models is inherently difficult [levin_energy_2023], mainly for two reasons:
@@
For one, while the literature has reached a broad consensus that VRE-based energy system models need to capture the spatio-temporal distribution of weather variables to find cost-effective and resilient system designs [collins_impacts_2018, schlachtberger_cost_2018, staffell_increasing_2018, gotske_designing_2024, goke_stabilized_2024, grochowicz_intersecting_2023], LDES has been shown to be particularly sensitive to the interannual variability in these variables [dowling_long-duration_2021, ruhnau_storage_2022, kittel_coping_2025], which is linked to the fact that LDES derives its system value from extreme system states happening rather infrequently. Recent literature contributions have sought to make the inclusion of a large number of weather years at a high granularity computationally feasible [grochowicz_intersecting_2023, goke_stabilized_2024, jacobson_computationally_2024]. However, most studies determine optimal LDES capacities based on a single weather year potentially leading to highly sub-optimal capacity choices when evaluated against other samples from the (unknown) weather distribution [zeyringer_designing_2018].
@@
For two, CEMs, even those taking into account interannual variability by including multiple weather years, commonly assume perfect foresight in the dispatch stage. This is another practical concession made to maintain computational tractability in light of high spatial, temporal and technological granularity commonly found in these models [conejo_generation_2016]. Assuming perfect foresight is particularly consequential for the operation of LDES. For example, LDES injection decisions in summer would be taken under perfect knowledge of the looming VRE scarcity events and necessary storage withdrawal in winter, which may lead to very unrealistic LDES dispatch patterns, LDES system value and consequently capacity decisions. In reality, a LDES or system operator would only have limited foresight based on (probabilistic) weather forecasts.
@@
Some studies have introduced heuristic decision rules to eliminate effectively foresight in LDES operations [dupre_la_tour_towards_2023, brown_price_2025], imposing to sell and buy energy at fixed prices or at prices as a function of the current storage level. These approaches are likely to underestimate LDES system value and may bias capacity decisions. By contrast, the literature on hydroelectric power systems has long dealt with optimal storage dispatch under inflow uncertainty [scott_modelling_1996, steeger_optimal_2014]. A fully renewable energy system based on wind and solar and relying on LDES for balancing is in many respects similar to a hydro-dominated system [blanchard_strategic_2024], where random water inflows are replaced by random feed-in of wind and solar irradiation.
@@
This paper has three main contributions: first, we present capacity expansion model for fully renewable, wind- and solar-based European energy systems subject to both demand-side (electrified heating) and supply-side (variable renewable energy) weather uncertainty. Second, we isolate the effect of limited foresight by comparing the multi-stage formulation to a two-stage formulation that takes into account interannual weather variability but assumes perfect foresight in the dispatch stage. We investigate differences in LDES operations and consequences for optimal capacity decisions. Third, we derive LDES bidding curves implicit in the optimal policy of the multi-stage formulation and discuss their dependence on the price of backup or loss of load. We also show that LDES under weather uncertainty has important implications for the price formation on fully renewable energy-only markets and thereby contribute to the current debate on the viability of energy-only markets without fuel costs [brown_price_2025, antweiler_new_2025].
@@ 2 Capacity expansion planning of renewable energy systems under weather uncertainty
The sensitivity of VRE-based energy systems to interannual weather variability and extreme weather events is well documented [pfenninger_dealing_2017, zeyringer_designing_2018, staffell_increasing_2018, bloomfield_quantifying_2021, craig_overcoming_2022, grochowicz_using_2024]. Consequently, it has become common practice to include more than one ’typical meteorological year’ in capacity expansion models [grochowicz_using_2024]. However, taking into account interannual variability does not necessarily account for weather uncertainty even if the terms are often conflated in this context. Weather uncertainty means that the realization of relevant weather variables in the future is not known to the decision maker by the time they make their decision and the problem becomes stochastic [conejo_decision_2010]. In other words, the decision maker has limited foresight with respect to random weather variables. A deterministic CEM not accounting for uncertainty has a single decision stage and the decision maker decides on capacity and dispatch at once. In a stochastic CEM information arrives and decisions are made in sequence.
@@
Panel A of Figure 1 represents the most common way in which the joint optimization of capacity and dispatch is repeated for every weather year. Resulting in a range of model outputs, this type can be considered a sensitivity analysis and not decision-making under weather uncertainty [levin_energy_2023]. In the stochastic programming literature this approach corresponds to a wait-and-see solution [shapiro_lectures_2009]. Some analyses have added a second step in which the system design based on one weather year is tested for dispatch on another year to find resilient system specifications among all output candidates [gotske_designing_2024, ruggles_planning_2024].
Panel B of the same figure represents contributions that solve the dispatch stage of a CEM over potentially long sequential periods of up to 70 weather years [brown_ultra-long-duration_2023]. Especially used in the context of LDES to capture interannual storage dynamics [zeyringer_designing_2018, dowling_role_2020, ruhnau_storage_2022], capacity decisions are made in view of many weather years but potentially distorted by extreme levels of perfect foresight. For example, Brown and Hampp [brown_ultra-long-duration_2023] show storage level trajectories that imply overarching discharge period, which define optimal energy capacity, of nearly a decade.
Panel C represents cases where each dispatch year is optimized in isolation but the capacity mix is constrained to be the same across weather years. In stochastic programming this corresponds to a two-stage stochastic program [shapiro_lectures_2009]. Recent literature contributions have sought to make models of this type computationally tractable [goke_stabilized_2024, grochowicz_intersecting_2023, jacobson_computationally_2024]. While they only differ from Panel B in that they do not consider dispatch in sequence, it represents a capacity decision under weather uncertainty, albeit with perfect foresight during dispatch.
Panel D shows a multi-stage stochastic capacity expansion model [shapiro_lectures_2009] abandoning the perfect foresight assumption during the dispatch stage. Existing literature contributions are mainly found for hydro-dominated systems, e.g. [rebennack_generation_2014, thome_stochastic_2019, hole_capacity_2025], some of which consider stochastic wind and solar generation next to random inflows. While Blanchard [blanchard_strategic_2024] and Hummelen et al. [hummelen_exploring_2024] explicitly consider hydrogen storage with limited foresight under solar and wind uncertainty, they only consider a partial capacity choice between electrolyzers and storage to hedge against hydrogen supply disruptions or use a very reduced scenario tree respectively.
@@
This work isolates the impact of perfect foresight on LDES operations and direct as well as indirect consequences for optimal capacity decisions in fully renewable wind- and solar-based energy systems by comparing the outcomes of model types D and C. The remainder of this section introduces formally a CEM of type D and discusses necessary assumptions and features of the resulting model. We use the terms system planner and system operator interchangeably.
Figure 1: Model setups to deal with weather variability and uncertainty: Panel A represents single year deterministic models. Panel B is a deterministic models that include a (long) sequence of weather years in the dispatch stage. Panel C is a two-stage stochastic setup in which the capacity decision is taken under uncertainty. There is perfect foresight in the dispatch stage. Lastly, Panel D represents the multi-stage stochastic setup that limits foresight in the dispatch stage.
@@ 2.1 Modeling capacity expansion with limited weather foresight
A CEM of type D leads to a multi-stage stochastic program (MSSP), where the first stage $t=0$ corresponds to the capacity decision and stages $t=1,\dots,T$ are dispatch stages. Each dispatch stage consists of $H(t)=|\mathcal{H}(t)|$ operational periods. When in dispatch stage $t$ the system operator has perfect foresight for all operational periods in the stage $t$, $\mathcal{H}(t)$. Weather realizations beyond $t$ are uncertain and the system operator needs to take the distribution of future outcomes into account when making a decision in stage $t$. Generally, this leads to a sequence of nested expectations, where each expectation is conditional on the past decisions. We provide a general problem formulation in Section SI.5.
The solution to such a MSSP is a policy, a set of functions that maps system states to capacity or dispatch choices. A system state is described by two components: 1) the exogenous weather realization and 2) an endogenous state resulting from previous decisions, such as capacity choices from stage 00 or the LDES storage level inherited from the previous dispatch stage.
@@
$\min_{\begin{subarray}{c}\{g_{r,t,h},f_{s,t,h},h_{s,T,h},g^{inf}_{T,h},e_{s,t,h}\}\\ \in\mathcal{O}_{t}(G_{r},F_{s},H_{s},E_{s},e_{s,t-1,H},\xi_{t})\end{subarray}}\sum_{h}\left[\sum_{r}o_{r}g_{r,t,h}+vg^{inf}_{t,h}\right]$ $+\mathbb{E}_{t}\left[{Q}_{t+1}(e_{s,t,H},G_{r},F_{s},H_{s},E_{s},e_{s}^{ini})\right]$
The optimal value function $Q_{t}(\cdot)$. is a function of state variables, where $G_{r},F_{s},H_{s},E_{s}$ are decisions for generation capacities, storage power and storage energy capacities from stage 00 respectively. $e_{s,t-1,H}$ is the storage ending level of storage $s$ in the previous stage $t-1$, $e_{s}^{ini}$ is the initial storage level that needs to be met again at the end of stage $T$ (see Section 2.3) and $\xi_{t}$ is the realization of the random weather vector $\xi_{t}\in\Xi_{t}$, which contains random capacity factors for wind and solar PV, hydro flows, heating demand and heat pump efficiencies for every period $h\in\mathcal{H}(t)$. These state variables affect the feasible space for the dispatch decisions in $t$, $\mathcal{O}_{t}(\cdot)$. The subproblem in $t$ is to minimize the sum of the stage’s dispatch costs, given by the sum of variable generation costs $\sum_{r}o_{r}g_{r,t,h}$ and the costs of load shedding $vg_{t,h}^{inf}$ and the expected future system costs, the cost-to-go. The latter are affected by the decisions in $t$ through the ending storage level $e_{s,t,H}$. In a subproblem the trade-off boils down to reducing the current stages dispatch costs at the expense of increasing the cost-to-go as the future is left with a lower storage level $e_{s,t,H}$. In stage $T$ the cost-to-go term is equal to 00. We introduce a terminal condition to avoid greedy storage discharging, which is discussed in Section 2.3.
@@
$\min_{\{G_{r},F_{s},H_{s},E_{s},e_{s}^{ini}\}\in\mathcal{C}}\sum_{r}c_{r}G_{r}+\sum_{s}\left(c_{s}^{f}F_{s}+c_{s}^{h}H_{s}+c_{s}^{e}E_{s}\right)+\mathbb{E}_{0}\left[Q_{1}(G_{r},F_{s},H_{s},E_{s},e_{s}^{ini})\right]$ (1b)
@@
Broadly, SDDP decomposes the nested problem by replacing the cost-to-go functions $\mathcal{Q}_{t+1}(\cdot):=\mathbb{E}_{t}[Q_{t+1}(\cdot)]=N^{-1}_{t}\sum_{i}Q_{t+1}(\cdot,\xi_{t+1,i})$ by cutting plane approximations $\hat{\mathcal{Q}}_{t+1}$, which are defined by a collection of linear constraints. Each subproblem is now a small linear program (LP) that is easy to solve on its own. SDDP improves the approximations $\{\hat{\mathcal{Q}}_{t}\}$ iteratively by sampling weather year sample paths from $\Xi$, solving each subproblem in sequence given the sampled weather year and given the current approximations of the cost-to-go functions by passing forward the resulting (endogenous) state variables (think capacities and storage levels). In a subsequent backward pass the algorithm collects feedback from the subproblems at these trial points, Benders cuts, that are added to the cost-to-go function approximations. We defer a detailed description of the algorithm to the Supplemental Information, Section SI.6.
@@
The perfect foresight benchmark model is of type C. We provide additional details in the Supplemental Information, Section SI.5. We refer to the two models as Limited Foresight (LF) and Perfect Foresight (PF) respectively.

### v2

@@ Abstract
Long-duration energy storage (LDES) is a key component for fully renewable, sector-coupled energy systems based on wind and solar. While capacity expansion planning has begun to take into account interannual weather variability, it often ignores weather uncertainty and limited foresight in capacity and operational decisions. We build a stochastic capacity expansion model for fully decarbonized energy systems with LDES in Europe accounting for weather uncertainty - isolating the effect of limited foresight by comparing it to a perfect foresight benchmark. Under limited foresight, LDES acts as a hedge against extreme system states operating defensively and exhibiting a stockpiling effect absent under perfect foresight. Solar PV gains in system value for its higher predictability with up to 25% higher capacities versus the benchmark while onshore wind capacities are lower. We shed light on the underlying mechanisms by deriving implicit LDES bidding curves. We show that LDES bids reflect the costs and the weather-dependent probability of extreme system states conditional on the current system state. This has important implications for the price formation on renewable electricity markets, as a wide and continuous range of probabilistic LDES bids alleviates concerns of extreme price disparity at high renewable shares.
@@ 1 Introduction
The energy sector is at the center of climate change mitigation efforts around the world [international_energy_agency_net_2024, intergovernmental_panel_on_climate_change_ipcc_technical_2023]. Variable renewable energy (VRE), including wind and solar energy, is likely to supply the majority of the energy in a decarbonized system for its ubiquitous availability, cost competitiveness and ease of deployment [schlachtberger_cost_2018, victoria_speed_2022, brown_response_2018, breyer_history_2022, luderer_impact_2022, hansen_status_2019]. The resulting system dependency on weather variables such as wind speeds, solar irradiation, water inflows and temperatures due to electrified heating and cooling demand creates flexibility requirements on various temporal and spatial scales [lund_review_2015, levin_energy_2023].
@@
Long-duration energy storage (LDES) is considered a key enabler of very high VRE shares in a decarbonized future by providing seasonal or interannual balancing and bridging periods of low solar and wind feed-in that may last multiple days or even weeks, so-called Dunkelflauten [lopez_prol_economics_2021, dowling_role_2020, ruhnau_storage_2022, sepulveda_design_2021, kittel_coping_2025, staadecker_value_2024]. LDES is not well defined but often used to refer to storage technologies that offer durations between ten and hundreds of hours and which are characterized by low energy capacity costs [dowling_long-duration_2021, albertus_long-duration_2020, brown_ultra-long-duration_2023]. While several technology options exist, most of the academic literature, planned legislation and current market developments in Europe and the US suggest that hydrogen storage in salt caverns in conjunction with electrolysis and hydrogen gas turbines (power-to-gas-to-power) will play a dominant role [dowling_role_2020, federal_ministry_for_economic_affairs_and_climate_action_of_germany_weisbuch_2025, rwe_erweiterungsprojekt_2024, uniper_press_2024].
@@
Capacity expansion models (CEM) are well-established and important tools for policymakers, investors and other stakeholders in the energy transition. Yet, the representation of LDES in these models is inherently difficult [levin_energy_2023], mainly for two reasons:
@@
First, while the literature has reached a broad consensus that VRE-based energy system models need to capture the spatio-temporal distribution of weather variables to find cost-effective and resilient system designs [collins_impacts_2018, schlachtberger_cost_2018, staffell_increasing_2018, gotske_designing_2024, goke_stabilized_2024, grochowicz_intersecting_2023], LDES is particularly sensitive to the interannual variability in these variables [dowling_long-duration_2021, ruhnau_storage_2022, kittel_coping_2025], which is linked to the fact that LDES derives its system value from extreme system states happening rather infrequently. Recent literature contributions have sought to make the inclusion of a large number of weather years at a high granularity computationally feasible [grochowicz_intersecting_2023, goke_stabilized_2024, jacobson_computationally_2024]. However, most studies determine optimal LDES capacities based on a single weather year, potentially leading to highly sub-optimal capacity choices when evaluated against other samples from the (unknown) weather distribution [zeyringer_designing_2018].
@@
Second, CEMs, even those taking into account interannual variability by including multiple weather years, commonly assume perfect foresight in the dispatch stage. This is another practical concession made to maintain computational tractability in light of high spatial, temporal and technological granularity commonly found in these models [conejo_generation_2016]. Assuming perfect foresight is particularly consequential for the operation of LDES. For example, LDES injection decisions in summer would be taken under perfect knowledge of the looming VRE scarcity events and necessary storage withdrawal in winter, which may lead to very unrealistic LDES dispatch patterns, LDES system value and consequently capacity decisions. In reality, a LDES or system operator would only have limited foresight based on (probabilistic) weather forecasts.
@@
Some studies have introduced heuristic decision rules to eliminate effectively foresight in LDES operations [dupre_la_tour_towards_2023, brown_price_2025]. These rules impose to sell and buy energy at fixed prices or determine bids as a function of the current storage level. These approaches are likely to underestimate LDES system value and may bias capacity decisions. By contrast, the literature on hydroelectric power systems has long dealt with optimal storage dispatch under inflow uncertainty [scott_modelling_1996, steeger_optimal_2014]. A fully renewable energy system based on wind and solar and relying on LDES for balancing is in many respects similar to a hydro-dominated system [blanchard_strategic_2024], where random water inflows are replaced by random feed-in of wind and solar irradiation.
@@
This paper has three main contributions: first, we present a capacity expansion model for fully renewable, wind- and solar-based European energy systems subject to both demand-side (electrified heating) and supply-side (variable renewable energy) weather uncertainty. Second, we isolate the effect of limited foresight by comparing the multi-stage formulation to a two-stage formulation that takes into account interannual weather variability but assumes perfect foresight in the dispatch stage. We investigate differences in LDES operations and consequences for optimal capacity decisions. Third, we derive LDES bidding curves implicit in the optimal policy of the multi-stage formulation and discuss their dependence on the cost of backup or loss of load. We also show that LDES bids under weather uncertainty have important implications for the price formation on fully renewable energy-only markets and thereby contribute to the current debate on the viability of energy-only markets without fuel costs [brown_price_2025, antweiler_new_2025].
@@ 2 Capacity expansion planning of renewable energy systems under weather uncertainty
The sensitivity of VRE-based energy systems to interannual weather variability and extreme weather events is well documented [pfenninger_dealing_2017, zeyringer_designing_2018, staffell_increasing_2018, bloomfield_quantifying_2021, craig_overcoming_2022, grochowicz_using_2024]. Consequently, it has become common practice to include more than one ’typical meteorological year’ in capacity expansion models [grochowicz_using_2024]. However, taking into account interannual variability does not necessarily account for weather uncertainty even if the terms are often conflated in this context. Weather uncertainty means that the realization of relevant weather variables in the future is not known to the decision maker by the time they make their decision and the problem becomes stochastic [conejo_decision_2010]. In other words, the decision maker has limited foresight with respect to random weather variables. A deterministic CEM, not accounting for uncertainty, has a single decision stage and the decision maker decides on capacity and dispatch at once. In a stochastic CEM, information arrives and decisions are made in sequence.
@@
Panel A of Figure 1 represents the most common way, in which the joint optimization of capacity and dispatch is repeated for every weather year. Resulting in a range of model outputs, this type can be considered a sensitivity analysis [levin_energy_2023]. In the stochastic programming literature, this approach corresponds to a wait-and-see solution [shapiro_lectures_2009]. Some analyses have added a second step in which the system design based on one weather year is tested for dispatch on another year to find resilient system specifications among all output candidates [gotske_designing_2024, ruggles_planning_2024].
[…]
Panel C represents cases where each dispatch year is optimized in isolation but the capacity mix is constrained to be the same across weather years. In stochastic programming, this corresponds to a two-stage stochastic program [shapiro_lectures_2009]. Recent literature contributions have sought to make models of this type computationally tractable [goke_stabilized_2024, grochowicz_intersecting_2023, jacobson_computationally_2024]. While they only differ from Panel B in that they do not consider dispatch in sequence, it represents a capacity decision under weather uncertainty, albeit with perfect foresight during dispatch.
Panel D shows a multi-stage stochastic capacity expansion model [shapiro_lectures_2009] abandoning the perfect foresight assumption during the dispatch stage. Existing literature contributions are mainly found for hydro-dominated systems, e.g. [rebennack_generation_2014, thome_stochastic_2019, hole_capacity_2025], some of which consider stochastic wind and solar generation next to random inflows. While Blanchard [blanchard_strategic_2024] and Hummelen et al. [hummelen_exploring_2024] explicitly consider hydrogen storage with limited foresight under solar and wind uncertainty, they only consider a partial capacity choice between electrolyzers and storage to hedge against hydrogen supply disruptions or use a very reduced scenario tree, respectively.
@@
This work isolates the impact of perfect foresight on LDES operations and the for optimal capacity decisions in fully renewable wind- and solar-based energy systems by comparing the outcomes of model types D and C. The remainder of this section introduces formally a CEM of type D and discusses necessary assumptions and features of the resulting model. We use the terms system planner and system operator interchangeably.
Figure 1: Model setups to deal with weather variability and uncertainty: Panel A represents single-year deterministic models. Panel B is a deterministic model that includes a (long) sequence of weather years in the dispatch stage. Panel C is a two-stage stochastic setup in which the capacity decision is taken under uncertainty. There is perfect foresight in the dispatch stage. Lastly, Panel D represents the multi-stage stochastic setup that limits foresight in the dispatch stage.
@@ 2.1 Modeling capacity expansion with limited weather foresight
A CEM of type D leads to a multi-stage stochastic program (MSSP), where the first stage $t=0$ corresponds to the capacity decision and stages $t=1,\dots,T$ are dispatch stages. Each dispatch stage consists of $H(t)=|\mathcal{H}(t)|$ operational periods. When in dispatch stage $t$, the system operator has perfect foresight for all operational periods $h\in\mathcal{H}(t)$ in the stage, $\mathcal{H}(t)$. Weather realizations beyond $t$ are uncertain and the system operator needs to take the distribution of future outcomes into account when making a decision in stage $t$. Generally, this leads to a sequence of nested expectations, where each expectation is conditional on the past decisions. We provide a general problem formulation in Section SI.5.
The solution to such an MSSP is a policy, a set of functions that maps system states to capacity or dispatch choices. A system state is described by two components: 1) the exogenous weather realization, and 2) an endogenous state resulting from previous decisions, such as capacity choices from stage 00 or the LDES storage level inherited from the previous dispatch stage.
@@
$\min_{\begin{subarray}{c}\{g_{r,t,h},f_{s,t,h},h_{s,t,h},g^{inf}_{T,h},e_{s,t,h}\}\\ \in\mathcal{O}_{t}(G_{r},F_{s},H_{s},E_{s},e_{s}^{ini},e_{s,t-1,H},\xi_{t})\end{subarray}}\sum_{h}\left[\sum_{r}o_{r}g_{r,t,h}+vg^{inf}_{t,h}\right]$ $+\mathbb{E}_{t}\left[{Q}_{t+1}(e_{s,t,H},G_{r},F_{s},H_{s},E_{s},e_{s}^{ini},\xi_{t+1})\right]$
The optimal value function $Q_{t}(\cdot)$ is a function of state variables, where $G_{r},F_{s},H_{s},E_{s}$ are decisions for generation capacities, storage power and storage energy capacities from stage 00 respectively. $e_{s,t-1,H}$ is the storage ending level of storage $s$ in the previous stage $t-1$, $e_{s}^{ini}$ is the initial storage level that needs to be met again at the end of stage $T$ (see Section 2.3) and $\xi_{t}\in\Xi_{t}$ is the random weather vector, which contains random capacity factors for wind and solar PV, hydro flows, heating demand and heat pump efficiencies for every period $h\in\mathcal{H}(t)$. These state variables affect the feasible space for the dispatch decisions in $t$, $\mathcal{O}_{t}(\cdot)$. The subproblem in $t$ is to minimize the sum of the stage’s dispatch costs, given by the sum of variable generation costs $\sum_{r}o_{r}g_{r,t,h}$ and the costs of load shedding $vg_{t,h}^{inf}$ in every period and the expected future system costs, the cost-to-go. The latter are affected by the decisions in $t$ through the ending storage level $e_{s,t,H}$. The trade-off boils down to reducing the current stage’s dispatch costs at the expense of increasing the cost-to-go as the future is left with a lower storage level $e_{s,t,H}$. In stage $T$ the cost-to-go term is equal to 00. We introduce a terminal condition to avoid greedy storage discharging, which is discussed in Section 2.3.
@@
$\min_{\{G_{r},F_{s},H_{s},E_{s},e_{s}^{ini}\}\in\mathcal{C}}\sum_{r}c_{r}G_{r}+\sum_{s}\left(c_{s}^{f}F_{s}+c_{s}^{h}H_{s}+c_{s}^{e}E_{s}\right)+\mathbb{E}_{0}\left[Q_{1}(G_{r},F_{s},H_{s},E_{s},e_{s}^{ini},\xi_{1})\right]$ (1b)
@@
Broadly, SDDP decomposes the nested problem by replacing the cost-to-go functions $\mathcal{Q}_{t+1}(\cdot):=\mathbb{E}_{t}[Q_{t+1}(\cdot)]=N^{-1}_{t}\sum_{i}Q_{t+1}(\cdot,\xi_{t+1,i})$ by cutting plane approximations $\hat{\mathcal{Q}}_{t+1}$, which are defined by a collection of linear constraints, so-called Benders cuts. Each subproblem is now a small linear program (LP) that is easy to solve on its own.
SDDP improves the approximations $\{\hat{\mathcal{Q}}_{t}\}$ iteratively using a forward and a backward pass. A forward pass begins with sampling a weather year path from $\Xi$. Next, each subproblem is solved in sequence given the sampled weather year and the current cost-to-go function approximations. The resulting endogenous state variables, capacities and storage levels, are passed from one subproblem to the next.
In a subsequent backward pass the algorithm moves through the stages in reverse order. Conditional on the incoming endogenous state variables defined in the forward pass, it collects feedback from each subproblem in the form of Benders cuts. These cuts improve the cost-to-go function approximation used in the preceding stage, respectively. We defer a detailed description of the algorithm to the Supplemental Information, Section SI.6.
@@
The perfect foresight benchmark model is of type C. We provide additional details in the Supplemental Information, Section SI.5. We refer to the two models as Limited Foresight (LF) and Perfect Foresight (PF), respectively.